If the difference between the largest and the smallest number from a set of five
numbers is 0, what is the difference between the second largest and the third largest
numbers?
1
Expert's answer
2020-04-20T18:12:11-0400

Answer:0

"Let\;a\geq b\geq c\geq d\geq e\;are\;our\;5\;numbers.\\a-e=0\Rightarrow a=e\\a\geq b\Rightarrow e\geq b\geq e\Rightarrow e=b\\a\geq c\Rightarrow e\geq c\geq e\Rightarrow e=c\\b-c=e-e=0"